2m=2-9n
21n=4-6m
solve for m and n.

Answer by chripchrip(6)

m=(2-9n)/2
substitue m to the equation 21n=4-6m
21n=4-6(2-9n)/2
21n*2=4*2-6(2-9n)
42n=8-12+54n
-12n=-4
n=1/3
so 2m=2-9(1/3)
2m=2-3
2m=-1
m=-1/2